Joseph wishes to have some flyers delivered for
his grocery business. Speedy Post charges a fee
of $200 plus 50 cents per flyer. Quick Send
charges a fee of $100 plus 80 cents per flyer. For
what number of flyers will the cost be the same
for either delivery company?

Answer by addingup(3677)

200 + 0.50x = 100 + 0.80x
100 = 0.30x
x = 100/0.30 = 333.33
The cost will be the same at approximately 333 flyers (you need whole flyers)